THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 80 — 1 April 2006

http://www.astro.keele.ac.uk/MCnews

Editors: Jacco van Loon and Snežana Stanimirović

Editorial

Dear Colleagues,

It is our pleasure to present the 80^{th} issue of the Magellanic Clouds Newsletter. We would like to express our thanks to the past editors of the Magellanic Clouds Newsletter over the last ten years, in particular You-Hua Chu who was one of its founders and who has been involved all the while. After a brief lapse, the Magellanic Clouds Newsletter has now been relaunched, from a new web site — http://www.astro.keele.ac.uk/MCnews

Although the main purpose of the newsletter remains unchanged, we wish to emphasize the opportunity to post announcements of new tools, databases or meetings, as well as job opportunities that help our younger colleagues establish a career in astronomy.

We are delighted to see so many contributions in the current issue, including 30 refereed journal papers and a recent PhD thesis. This proves that the Magellanic Clouds community is a vibrant one. We hope that the rich diversity of topics forms an inspiration for further Magellanic research, and that the newsletter can serve as a platform to disseminate results and announcements.

The plan is for the Magellanic Clouds Newsletter to appear bimonthly. In the future, notification of issues of the newsletter will only be e-mailed to registered addresses. If you would like to continue to receive such notifications, please make sure that you have registered with the mailing list. This can be done on the web site. Do not forget to tell your students and colleagues about it, and let us know if you have any comments about the newsletter.

The next issue will be distributed on the 2^{nd} of June; the deadline for contributions is the 1^{st} of June.

Editorially Yours,

Jacco van Loon and Snežana Stanimirović

Refereed Journal Papers

The magnetic field of the Large Magellanic Cloud revealed through Faraday rotation

B. M. Gaensler^{1,2}, M. Haverkorn^{1}, L. Staveley-Smith^{3}, J. M. Dickey^{4}, N. M. McClure-Griffiths^{3}, J. R. Dickel^{5} and M. Wolleben^{6}

1Harvard-Smithsonian Center for Astrophysics, USA 2University of Sydney, Australia 3Australia Telescope National Facility, Australia 4University of Tasmania, Australia 5University of Illinois, USA 6Max-Planck Institut für Radioastronomie, Germany

We have measured the Faraday rotation toward a large sample of polarized radio sources behind the Large Magellanic Cloud (LMC), to determine the structure of this galaxy’s magnetic field. The magnetic field of the LMC consists of a coherent axisymmetric spiral of field strength ∼ 1 microgauss. Strong fluctuations in the magnetic field are also seen, on small (< 0.5 parsecs) and large (∼ 100 parsecs) scales. The significant bursts of recent star formation and supernova activity in the LMC argue against standard dynamo theory, adding to the growing evidence for rapid field amplification in galaxies.

Published in Science, v307, pp1610-1612 (2005) Available from astro-ph/0503226

Chemical analysis of carbon stars in the Local Group: I. The Small Magellanic Cloud and the Sagittarius dwarf spheroidal galaxy

P. de Laverny^{1}, C. Abia^{2}, I. Domínguez^{2}, B. Plez^{3}, O. Straniero^{4}, R. Wahlin^{5}, K. Eriksson^{5} and U.G. Jørgensen^{6}

1Observatoire de la Côte d’Azur, Dpt. Cassiopée UMR6202, 06304, Nice cedex 4, France 2Dpto. Física Teórica y del Cosmos, Universidad de Granada, E-18 071 Granada, Spain 3GRAAL, UMR5024, Université de Montpellier II, 34 095 Montpellier cedex 5, France 4INAF-Osservatorio di Collurania, 64 100 Teramo, Italy 5Department of Astronomy and Space Physics, Box 515, 75120 Uppsala, Sweden 6Niels Bohr Institute, Astronomical Observatory, Juliane Maries vej 30, 2 100 Copenhagen, Denmark

We present the first results of our ongoing chemical study of carbon stars in the Local Group of galaxies. We used spectra obtained with UVES at the 8.2 m Kueyen-VLT telescope and a new grid of spherical model atmospheres for cool carbon-rich stars which include polyatomic opacities, to perform a full chemical analysis of one carbon star, BMB-B 30, in the Small Magellanic Cloud (SMC) and two, IGI95-C1 and IGI95-C3, in the Sagittarius Dwarf Spheroidal (Sgr dSph) galaxy. Our main interest is to test the dependence on the stellar metallicity of the s-process nucleosynthesis and mixing mechanism occurring in AGB stars. For these three stars, we find important s-element enhancements with respect to the mean metallicity ([M/H]), namely [s/M]≈+1.0, similar to the figure found in galactic AGB stars of similar metallicity. The abundance ratios derived between elements belonging to the first and second s-process abundance peaks, corresponding to nuclei with a magic number of neutrons N = 50 (^{88}Sr, ^{89}Y, ^{90}Zr) and N = 82 (^{138}Ba, ^{139}La, ^{140}Ce, ^{141}Pr), agree remarkably well with the theoretical predictions of low mass (M < 3 M_{⊙}) metalpoor AGB nucleosynthesis models where the main source of neutrons is the ^{13}C(α, n)^{16}O reaction activated during the long interpulse phase, in a small pocket located within the He-rich intershell. The derived C/O and ^{12}C/^{13}C ratios are, however, more difficult to reconcile with theoretical expectations. Possible explanations, like the extrinsic origin of the composition of these carbon stars or the operation of a non-standard mixing process during the AGB phase (as the cool bottom process), are discussed on the basis of the collected observational constraints.

Published in A&A 446, 1107 (2006)

Echoes from Ancient Supernovae in the Large Magellanic Cloud

A. Rest^{1}, N. B. Suntzeff^{1}, K. Olsen^{1}, J.L. Prieto^{2}, R.C. Smith^{1}, D.L. Welch^{3}, A. Becker^{4}, M. Bergmann^{5}, A. Clocchiatti^{6}, K. Cook^{7}, A. Garg^{8}, M. Huber^{7}, G. Miknaitis^{4}, D. Minniti^{6}, S. Nikolaev^{7} and C. Stubbs^{8}

1CTIO/NOAO, Chile 2Ohio State University, USA 3McMaster University, Canada 4University of Washington, USA 5Gemini Observatory, Chile 6Pontifica Universidad Católica de Chile, Chile 7Lawrence Livermore National Laboratory, USA 8Harvard University, USA

In principle, the light from historical supernovae could still be visible as scattered-light echoes even centuries later. However, while echoes have been discovered around some nearby extragalactic supernovae well after the explosion, targeted searches have not recovered any echoes in the regions of historical Galactic supernovae. The discovery of echoes can allow us to pinpoint the supernova event both in position and age and, most importantly, allow us to acquire spectra of the echo light to type the supernova centuries after the direct light from the explosion first reached the Earth. Here we report on the discovery of three faint new variable surface brightness complexes with high apparent proper motion pointing back to well-defined positions in the Large Magellanic Cloud (LMC). These positions correspond to three of the six smallest (and likely youngest) previously catalogued supernova remnants, and are believed to be due to thermonuclear (Type Ia) supernovae. Using the distance and proper motions of these echo arcs, we estimate ages of 610 and 410 yr for the echoes #2 and #3.

Published in Nature, 428, 1132 (2005) Available from astro-ph/0510738 and from www.noao.edu/outreach/press/pr05/pr0510.html

The Initial Mass Function toward the Low-Mass End in the Large Magellanic Cloud with Hubble Space Telescope WFPC2 Observations

Gouliermis, D.^{1}, Brandner, W.^{1} and Henning, Th.^{1}

1Max-Planck-Institut für Astronomie, Heidelberg, Germany

We present V- and I-equivalent Hubble Space Telescope (HST) WFPC2 photometry of two areas in the Large Magellanic Cloud: the southern part of the stellar association LH 52, located on the western edge of the supershell LMC 4, and a field between two associations, which is located on the southwestern edge of the shell and accounts for the general background field of the galaxy. The HST WFPC2 observations reach magnitudes as faint as V=25 mag, much deeper than have been observed earlier in stellar associations in the LMC. We determine the mass function (MF) for main-sequence stars in the areas. Its slope in both areas is steeper for stars with masses M < 2 M_{⊙} (−4 < Γ < −6) than for stars of M > 2 M_{⊙} (−1 < Γ < −2). Thus, as far as the field of the LMC is concerned, the MF does not have a uniform slope throughout its observed mass range. The MF of the general field of the LMC was found previously to be steeper than the MF of a stellar association for massive stars with M > 5 M_{⊙}. We conclude that this seems to also be the case toward lower masses down to M ∼ 1 M_{⊙}. Our data allow us to construct the field-subtracted, incompleteness-corrected, main-sequence MF of the southwestern part of the young stellar association LH 52, which accounts for the initial mass function (IMF) of the system. Its mean slope is found to be comparable to, but more shallow than, a typical Salpeter IMF (Γ = −1.12 ± 0.24) for masses down to ∼ 1 M_{⊙}. We found indications that the IMF of the association probably is “top heavy,” owing to the large number of intermediate-mass stars in the field of the system, while the general LMC field is found to be responsible for the low-mass population, with M < 2 M_{⊙}, observed in both fields. This finding suggests that the local conditions seem to favor the formation of higher mass stars in associations, and not in the background field. No evidence for flattening of the IMF toward the low-mass regime or for a lower mass cutoff in the IMF was detected in our data.

Published in The Astrophysical Journal, Volume 623, Issue 2, pp. 846-859. Available from astro-ph/0411448

The Low-Mass Pre-Main-Sequence Population of the Stellar Association

LH 52 in the Large Magellanic Cloud Discovered with Hubble Space Telescope WFPC2 Observations

Gouliermis, D.^{1}, Brandner, W.^{1} and Henning, Th.^{1}

1Max-Planck-Institut für Astronomie, Heidelberg, Germany

We report on the serendipitous discovery of ∼500 low-mass candidate pre-main-sequence (PMS) stars in the vicinity of the stellar association LH 52 in the Large Magellanic Cloud. We present evidence that the red faint sequence of these stars seen in the CMD of LH 52 from Hubble Space Telescope (HST) WFPC2 observations belongs only to the association and follows almost perfectly isochrone models for PMS stars of masses down to ∼0.3 M_{⊙}. We find that this feature has a Galactic counterpart and that the mass spectrum of the candidate PMS stars in LH 52 seems to correspond to a Salpeter initial mass function with a slope Γ ∼ −1.26 in the mass range 0.8–1.4 M_{⊙}.

Published in The Astrophysical Journal, Volume 636, Issue 2, pp. L133-L136. Available from astro-ph/0512094

The Low-Mass Initial Mass Function of the field population in the Large Magellanic Cloud with Hubble Space Telescope WFPC2 Observations

Gouliermis, D.^{1}, Brandner, W.^{1} and Henning, Th.^{1}

1Max-Planck-Institut für Astronomie, Heidelberg, Germany

We present V- and I-equivalent HST/WFPC2 stellar photometry of an area in the Large Magellanic Cloud (LMC), located to the west of the bar of the galaxy, which accounts for the general background field of its inner disk. The WFPC2 observations reach magnitudes as faint as V=25 mag, and the large sample of more than 80,000 stars allows us to determine in detail the Present-Day Mass Function (PDMF) of the detected main-sequence stars, which is identical to the Initial Mass Function (IMF) for masses M< ∼1 M_{⊙}. The low-mass main-sequence mass function of the LMC field is found not to have a uniform slope throughout the observed mass range, i.e. the slope does not follow a single power law. This slope changes at about 1 M_{⊙} to become more shallow for stars with smaller masses down to the lowest observed mass of ∼ 0.7 M_{⊙}, giving clear indications of flattening for even smaller masses. We verified statistically that for stars with M< ∼1 M_{⊙} the IMF has a slope Γ around −2, with an indicative slope Γ ∼ −1.4 for 0.7< ∼M/M_{⊙}< ∼0.9, while for more massive stars the main-sequence mass function becomes much steeper with Γ ∼ −5. The main-sequence luminosity function (LF) of the observed field is in very good agreement with the Galactic LF as it was previously found. Taking into account several assumptions concerning evolutionary effects, which should have changed through time the stellar content of the observed field, we reconstruct qualitatively its IMF for the whole observed mass range (0.7< ∼M/M_{⊙}< ∼2.3) and we find that the number of observed evolved stars is not large enough to have affected significantly the form of the IMF, which thus is found almost identical to the observed PDMF.

Accepted for publication in The Astrophysical Journal Available from astro-ph/0508686

The Large Magellanic Cloud: diffuse interstellar bands, atomic lines and the local environmental conditions

N.L.J. Cox^{1}, M.A. Cordiner^{2}, J. Cami^{3}, B.H. Foing^{4}, L. Kaper^{1} and P. Ehrenfreund^{5}

1Astronomical Institute “Anton Pannekoek”, University of Amsterdam, Kruislaan 403, 1098 SJ, Amsterdam, Netherlands 2School of Chemistry, The University of Nottingham, University Park, Nottingham, NG7 2RD, United Kingdom 3Space Science Division, NASA Ames Research Center, Mail Stop 245-6, Moffett Field, California 94035, USA 4ESA/SCI-SR, ESTEC, P.O Box 299, 2200 AG Noordwijk, The Netherlands 5Astronomy Institute, Radboud University Nijmegen, P.O. Box 9010, 6500 GL Nijmegen, The Netherlands

The Large Magellanic Cloud (LMC) offers a unique laboratory to study the diffuse interstellar bands (DIBs) under conditions that are profoundly different from those in the Galaxy. DIB carrier abundances depend on several environmental factors, in particular the local UV radiation field. In this paper we present measurements of twelve DIBs detected in five lines of sight to early-type stars in the LMC, including the 30 Doradus region. From the high resolution spectra obtained with VLT/UVES we also derive environmental parameters that characterise the local interstellar medium (ISM) in the probed LMC clouds. These include the column density components (including total column density) for the atomic resonance lines of Na i, Ca ii, Ti ii, K i. In addition, we derive the H i column density from 21 cm line profiles, the total-to-selective visual extinction R_{V} and the gas-to-dust ratio N(H i)/A_{V}. Furthermore, from atomic line ratios we derive the ionisation balance and relative UV field strength in these environments. We discuss the properties of the LMC ISM in the context of DIB carrier formation. The behaviour of DIBs in the LMC is compared to that of DIBs in different local environmental conditions in the Milky Way. A key result is that in most cases the diffuse band strengths are weak (up to factor 5) with respect to Galactic lines of sight of comparable reddening, E_{B−V}. In the line of sight towards Sk-69 223 the 5780 and 5797 Å DIBs are very similar in strengths and profile to those observed towards HD 144217, which is typical of an environment exposed to a strong UV field. From the velocity analysis we find that DIB carriers (towards Sk-69 243) are better correlated with the ionised species like Ca ii than with neutrals (like Na i and CO). The most significant parameter that governs the behaviour of the DIB carrier is the strength of the UV field.

Published in 2006 Astronomy and Astrophysics 447 991

The Discovery of a Population of Pre-Main Sequence Stars in N66/NGC346 from Deep HST/ACS Images

A. Nota^{1,2}, M. Sirianni^{1,2}, E. Sabbi^{1}, M. Tosi^{3}, M. Clampin^{4}, J. Gallagher^{5}, M. Meixner^{1}, M.S. Oey^{6}, A. Pasquali^{7}, L.J. Smith^{8}, R. Walterbos^{9} and J. Mack^{1}

1Space Telescope Science Institute, 3700 San Martin Drive, Baltimore 2ESA, Space Telescope Operations Division 3INAF-Osservatorio Astronomico di Bologna, via Ranzani 1, Bologna, I-40127, Italy 4Goddard Space Flight Center, Greenbelt, Maryland, MD 20771, USA 5Department of Astronomy, University of Wisconsin, 475 North Charter Street, Madison, WI 53706-1582, USA 6Department of Astronomy, University of Michigan, 830 Dennison Buidling, Ann Arbor, Michigan, MI 48109, USA 7Max-Planck-Institut für Astronomie, Königstuhl 17, D-69117 Heidelberg, Germany 8Department of Physics and Astronomy, University College London, Gower Street, London, WC1E 6BT, GB 9Department of Astronomy, New Mexico State University, Department 4500, P.O. Box 30001, Las Cruces, NM 88003, USA

We report the discovery of a rich population of low mass stars in the young, massive star forming region N66/NGC346 in the Small Magellanic Cloud, from deep V, I and Hα images taken with the HST/ACS. These stars have likely formed together with the NGC346 cluster, ≃ 3–5 Myr ago. Their magnitude and colors are those of pre-main sequence stars in the mass range 0.6-3 M_{⊙}, mostly concentrated in the main cluster, but with secondary subclusters spread over a region across ∼ 45 pc. These subclusters appear to be spatially coincident with previously known knots of molecular gas identified in ground based and ISO observations. We show that N66/NGC346 is a complex region, being shaped by its massive stars, and the observations presented here represent a key step towards the understanding of how star formation occurred and has progressed in this low metallicity environment.

Accepted for publication in ApJ Letters Available from astro-ph/0602218 and from ST ScI Eprint #1713

CHANDRA ACIS Spectroscopy of N157B — A Young Composite Supernova Remnant in a Superbubble

Yang Chen^{1}, Q. Daniel Wang^{2,3}, Eric. V. Gotthelf^{4}, Bing Jiang^{1}, You-Hua Chu^{5} and Robert Gruendl^{5}

1Department of Astronomy, Nanjing University 2Department of Astronomy, University of Massachusetts at Amherst 3Institute for Advanced Study, Princeton University 4Columbia Astrophysics Laboratory, Columbia University 5Astronomy Department, University of Illinois at Urbana-Champaign

We present Chandra ACIS observations of N157B, a young supernova remnant (SNR) located in the 30 Doradus starformation region of the Large Magellanic Cloud. This remnant contains the most energetic pulsar known (Ė = 4.8×10^{38} erg s^{−1} cm^{−2}), which is surrounded by a bright nonthermal nebula that likely represents a toroidal pulsar wind terminal shock observed edge-on. We confirm the non-thermal nature of the comet-shaped X-ray emission feature and show that the spectral steepening of this feature away from the pulsar is quantitatively consistent with synchrotron cooling of shocked pulsar wind particles flowing downstream at a bulk velocity close to the speed of light. Around the cometary nebula we unambiguously detect a thermal component, which accounts for about 1/3 of the total 0.5 – 10 keV flux from the remnant. This thermal component is distributed among various clumps of metal-enriched plasma embedded in the low surface brightness X-ray-emitting diffuse gas. The relative metal enrichment pattern suggests that the mass of the supernova progenitor is > ∼20 M_{⊙}. A comparison of the X-ray data with Hubble Space Telescope optical images suggests that the explosion site is close to a dense cloud, against which a reflection shock is launched. The interaction between this reflection shock and the nebula has likely produced both its cometary shape and the surrounding thermal emission enhancement. SNR N157B is apparently expanding into the hot low-density interior of a nearby superbubble formed by the young OB association LH99, as revealed by Spitzer mid-infrared images. This scenario naturally explains the exceptionally large sizes of both the thermal and nonthermal components as well as the lack of an outer shell of the SNR. However, the real situation in the region is likely to be more complicated. We find that a round-shaped soft X-ray-emitting clump with distinct spectral properties may result from a separate oxygen-rich remnant. These results provide a rare glimpse into the SNR structure and evolution of a recent star-formation region.

Submitted to ApJ Available from astro-ph/0603123

Evolution of the Reverse Shock Emission from SNR 1987A

Kevin Heng^{1}, Richard McCray^{1}, Svetozar A. Zhekov^{2}, Peter M. Challis^{3}, Roger A. Chevalier^{4}, Arlin P.S. Crotts^{5}, Claes Fransson^{6}, Peter Garnavich^{7}, Robert P. Kirshner^{3}, Stephen S. Lawrence^{8}, Peter Lundqvist^{6}, Nino Panagia^{9},

C.S.J. Pun^{10}, Nathan Smith^{11}, Jesper Sollerman^{12} and Lifan Wang^{13}

1JILA, University of Colorado, 440 UCB, Boulder, CO 2Space Research Institute, Moskovska str. 6, Sofia-1000, Bulgaria 3Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138 4Department of Astronomy, University of Virginia, P.O. Box 3818, Charlottesville, VA 22903-0818 5Columbia Astrophysics Laboratory, Columbia University, 550 West 120th Street, New York, NY 10027 6Stockholm Observatory, Department of Astronomy, AlbaNova, SE-106 91, Stockholm, Sweden 7Department of Physics, University of Notre Dame, 225 Nieuwland Science Hall, Notre Dame, IN 46556 8Department of Physics and Astronomy, 151 Hofstra University, Hempstead, NY 11590 9Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218 10Department of Physics, University of Hong Kong, Pokfulam Road, Hong Kong, China 11Center for Astrophysics and Space Astronomy, University of Colorado, 389 UCB, Boulder, CO 80309 12Dark Cosmology Center, Niels Bohr Institute, University of Copenhagen, Juliane Maries Vej 30, DK-2100 Copenhagen Ø, Denmark 13Institute for Nuclear and Particle Astrophysics, E.O. Lawrence Berkeley National Laboratory, Berkeley, CA 94720

We present new (2004 July) G750L and G140L Space Telescope Imaging Spectrograph (STIS) data of the Hα and Lyα emission from supernova remnant (SNR) 1987A. With the aid of earlier data, from Oct 1997 to Oct 2002, we track the local evolution of Lyα emission and both the local and global evolution of Hα emission. We find that the average Hα intensity has increased locally by a factor of ∼ 3 for both blueshifted and redshifted emission, over periods of about 5 and 7 years, respectively. The average Lyα intensity has increased by a factor ∼ 9, over about 5 years, for both components. The most recent observations allow us, for the first time, to directly compare the Hα and Lyα emission from the same slit position and at the same epoch. Consequently, we find clear evidence that, unlike Hα, Lyα is reflected from the debris by resonant scattering. In addition to emission which we can clearly attribute to the surface of the reverse shock, we also measure comparable emission, in both Hα and Lyα, which appears to emerge from supernova debris interior to the surface. New observations taken through slits positioned slightly eastward and westward of a central slit show a departure from cylindrical symmetry in the Hα surface emission. No obvious asymmetry is seen in the interior emission. Using a combination of old and new observations, we construct a light curve of the total Hα flux, F, from the reverse shock, which has increased by a factor ∼ 4 over about 8 years. However, due to large systematic uncertainties, we are unable to discern between the two limiting behaviours of the flux − F ∝ t (self-similar expansion) and F ∝ t^{5} (halting of the reverse shock). Such a determination is important for constraining the rate of hydrogen atoms crossing the shock, which is relevant to the question of whether the reverse shock emission will vanish in < ∼7 years (Smith et al. 2005). Future deep, low- or moderate-resolution spectra are essential for accomplishing this task.

Accepted for publication in ApJ Available from astro-ph/0603151

NLTE models of line-driven stellar winds II. O stars in SMC

Jiri Krticka^{1}

1Faculty of Science, Masaryk University, Kotlarska 2, 611 37 Brno, Czech Republic

We calculate NLTE line-driven wind models of selected O stars in the spectral range of O4 to O9 in the Small Magellanic Cloud (SMC). We compare predicted basic wind properties, i.e. the terminal velocity and the mass-loss rate with values derived from observation. We found relatively good agreement between theoretical and observed terminal velocities. On the other hand, predicted mass-loss rates and mass-loss rates derived from observation are in a good agreement only for higher mass-loss rates. Theoretical mass-loss rates lower than approximately 10^{−7} M_{⊙} year^{−1} are significantly higher than those derived from observation. These results confirm the previously reported problem of weak winds, since our calculated mass-loss rates are in a fair agreement with predictions of Vink et al. (2001). We study multicomponent models for these winds. For this purpose we develop a more detailed description of wind decoupling. We show that the instability connected with the decoupling of individual wind elements may occur for low-density winds. In the case of winds with very low observed mass-loss rate the multicomponent effects are important for the wind structure, however this is not able to consistently explain the difference between predicted mass-loss rate and mass-loss rate derived from observation for these stars. Similar to previous studies, we found the dependence of wind parameters on the metallicity. We conclude that the wind mass-loss rate significantly increases with metallicity as Ṁ ∼ Z^{0.67}, whereas the wind terminal velocity on average depends on metallicity only slightly, namely v_{∞} ∼ Z^{0.06} (for studied stars).

Accepted for publication in MNRAS Available from astro-ph/0601129

The proper motion of the Large Magellanic Cloud: a reanalysis

Mario H. Pedreros^{1}, Edgardo Costa^{2} and René A. Méndez^{2}

1Departamento de Física, Facultad de Ciencias, Universidad de Tarapacá, Arica, Chile 2Departamento de Astronomía, Universidad de Chile, Santiago, Chile

We have determined the proper motion (PM) of the Large Magellanic Cloud (LMC) relative to four background quasi-stellar objects, combining data from two previous studies made by our group, and new observations carried out in three epochs not included the original investigations. The new observations provided a significant increase in the time base and in the number of frames, relative to what was available in our previous studies. We have derived a total LMC PM of μ = (+2.0±0.1) mas yr^{−1}, with a position angle of θ = (62.4±3.1)^{∘}. Our new values agree well with most results obtained by other authors, and we believe we have clarified the large discrepancy between previous results from our group. Using published values of the radial velocity for the center of the LMC, in combination with the transverse velocity vector derived from our measured PM, we have calculated the absolute space velocity of the LMC. This value, along with some assumptions regarding the mass distribution of the Galaxy, has in turn been used to calculate the mass of the Milky Way. Our measured PM also indicates that the LMC is not a member of a proposed stream of galaxies with similar orbits around our galaxy.

Published in Astronomical Journal Vol. 131, p. 1461

A study of the B and Be star population in the field of the LMC open cluster NGC2004 with VLT-FLAMES

C. Martayan^{1}, A.M. Hubert^{1}, M. Floquet^{1}, J. Fabregat^{2}, Y. Frémat^{3}, C. Neiner^{4,1}, P. Stee ^{5} and J. Zorec^{6}

1GEPI, UMR8111 du CNRS, Observatoire de Paris-Meudon, 92195 Meudon Cedex, France 2Universidad de Valencía, Instituto de Ciencia de los Materiales, P.O. Box 22085, 46071 Valencía, Spain 3Royal Observatory of Belgium, 3 Avenue Circulaire, 31180 Brussels, Belgium 4Instituut voor Sterrenkunde, KU Leuven, Celestijnenlaan 220B, 3001 Belgium 5Observatoire de la Côte d’Azur, avenue Nicolas Copernic, 06130 Grasse, France 6Institut d’Astrophysique de Paris (IAP), 98bis boulevard Arago, 75014 Paris, France

Observations of hot stars belonging to the young cluster LMC-NGC2004 and its surrounding region have been obtained with the VLT-GIRAFFE facilities in MEDUSA mode. 25 Be stars were discovered; the proportion of Be stars compared to B-type stars is found to be of the same order in the LMC and in the Galaxy fields. 23 hot stars were discovered as spectroscopic binaries (SB1 and SB2), 5 of these are found to be eclipsing systems from the MACHO database, with periods of a few days. About 75% of the spectra in our sample are polluted by hydrogen (Hα and Hγ), [S ii] and [N ii] nebular lines. These lines are typical of H ii regions. They could be associated with patchy nebulosities with a bi-modal distribution in radial velocity, with higher values (+335 km s^{−1}) preferentially seen inside the southern part of the known bubble LMC4 observed in H i at 21 cm.

Published in A&A, 445, 931 (2006)

Photometry of Magellanic Cloud clusters with the Advanced Camera for Surveys - II. The unique LMC cluster ESO 121-SC03

A. D. Mackey^{1}, M. J. Payne^{1} and G. F. Gilmore^{1}

1Institute of Astronomy, Cambridge, UK

We present the results of photometric measurements from images of the LMC cluster ESO 121-SC03 taken with the Advanced Camera for Surveys on the Hubble Space Telescope. Our resulting colour-magnitude diagram reaches 3 magnitudes below the main-sequence turn-off, and represents by far the deepest observation of this cluster to date. We also present similar photometry from ACS imaging of the accreted Sagittarius dSph cluster Palomar 12, used in this work as a comparison cluster. From analysis of its CMD, we obtain estimates for the metallicity and reddening of ESO 121-SC03: [Fe/H] = −0.97 ± 0.10 and E(V − I) = 0.04 ± 0.02, in excellent agreement with previous studies. The observed horizontal branch level in ESO 121-SC03 suggests this cluster may lie 20 per cent closer to us than does the centre of the LMC. ESO 121-SC03 also possesses a significant population of blue stragglers, which we briefly discuss. Our new photometry allows us to undertake a detailed study of the age of ESO 121-SC03 relative to Pal.

12 and the Galactic globular cluster 47 Tuc. We employ both vertical and horizontal differential indicators on the CMD, calibrated against isochrones from the Victoria-Regina stellar models. These models allow us to account for the different α-element abundances in Pal. 12 and 47 Tuc, as well as the unknown run of α-elements in ESO 121-SC03. Taking a straight error-weighted mean of our set of age measurements yields ESO 121-SC03 to be 73 ± 4 per cent the age of 47 Tuc, and 91 ± 5 per cent the age of Palomar 12. Pal. 12 is 79 ± 6 per cent as old as 47 Tuc, consistent with previous work. Our result corresponds to an absolute age for ESO 121-SC03 in the range 8.3 − 9.8 Gyr, depending on the age assumed for 47 Tuc, therefore confirming ESO 121-SC03 as the only known cluster to lie squarely within the LMC age gap. We briefly discuss a suggestion from earlier work that ESO 121-SC03 may have been accreted into the LMC system.

Accepted for publication in MNRAS Available from astro-ph/0603472

Effects of metallicity, star-formation conditions, and evolution in B and Be stars. I: Large Magellanic Cloud, field of NGC 2004.

C. Martayan^{1}, Y. Frémat^{2}, A.-M. Hubert^{1}, M. Floquet^{1}, J. Zorec ^{3} and C. Neiner^{1}

1GEPI, UMR 8111 du CNRS, Observatoire de Paris-Meudon, 92195 Meudon Cedex, France 2Royal Observatory of Belgium, 3 avenue circulaire, 1180 Brussels, Belgium 3Institut d’Astrophysique de Paris (IAP), 98bis boulevard Arago, 75014 Paris, France

To statistically study the effects of the metallicity, star-formation conditions, and evolution on the behaviour of massive stars and, more particularly, of B and Be stars, we observed large samples of stars in the Magellanic Clouds for the first time. In this article we present the first part of this study. Spectroscopic observations of hot stars belonging to the young cluster LMC-NGC 2004 and its surrounding region were carried out with the VLT-GIRAFFE facilities in MEDUSA mode. We determined the fundamental parameters (T_{eff}, log g, v sin i, and radial velocity) for all B and Be stars in the sample thanks to a code developed in our group. The effect of fast rotation (stellar flattening and gravitational darkening) are taken into account in this study. We also determined the age of observed clusters. We then compared the mean v sin i obtained for field and cluster B and Be stars in the Large Magellanic Cloud (LMC) with the ones in the Milky Way (MW). We find, in particular, that Be stars rotate faster in the LMC than in the MW, in the field as well as in clusters. We discuss the relations between v sin i, metallicity, star-formation conditions, and stellar evolution by comparing the LMC with the MW. We conclude that Be stars began their main sequence life with an initial rotational velocity higher than the one for B stars. It is probable that only part of the B stars, those with a sufficient initial rotational velocity, can become Be stars. This result may explain the differences in the proportion of Be stars in clusters with similar ages.

Accepted for publication in A&A Available from astro-ph/0601240

Integrated-light VRI imaging photometry of globular clusters in the Magellanic Clouds

Paul Goudfrooij^{1}, Diane Gilmore^{1}, Markus Kissler-Patig^{2} and Claudia Maraston^{3}

1STScI 2ESO 3Astrophysics, University of Oxford

We present accurate integrated-light photometry in Johnson/Cousins V, R, and I for a sample of 28 globular clusters in the Magellanic Clouds. The majority of the clusters in our sample have reliable age and metallicity estimates available in the literature. The sample encompasses ages between 50 Myr and 7 Gyr, and metallicities ([Fe/H]) between −1.5 and 0.0 dex. The sample is dominated by clusters of ages between roughly 0.5 and 2 Gyr, an age range during which the bolometric luminosity of simple stellar populations is dominated by evolved red giant branch stars and thermally pulsing asymptotic giant branch (TP-AGB) stars whose theoretical colours are rather uncertain. The VRI colours presented in this paper have been used to calibrate stellar population synthesis model predictions.

Accepted for publication in MNRAS Available from astro-ph/0603355

AGB stars in the Magellanic Clouds. II. The rate of star formation across the LMC

M.-R.L. Cioni^{1}, L. Girardi^{2}, P. Marigo^{3} and H.J. Habing^{4}

1SUPA, School of Physics, University of Edinburgh, IfA, Blackford Hill, Edinburgh EH9 3HJ, UK 2Osservatorio Astronomico di Trieste, INAF, via G. B. Tiepolo 11, 34131 Trieste, Italy 3Dipartimento di Astronomia, Universit di Padova, Vicolo dell’Osservatorio 2, 35122 Padova, Italy 4Sterrewacht Leiden, Niels Bohrweg 2, 2333 RA Leiden, The Netherlands

This article compares the distribution of Ks magnitudes of Large Magellanic Cloud (LMC) asymptotic giant branch (AGB) stars obtained from the DENIS and 2MASS data with theoretical distributions. These have been constructed using up-to-date stellar evolution calculations for low and intermediate-mass stars, and in particular for thermally pulsing AGB stars. A fit of the magnitude distribution of both carbon- and oxygen-rich AGB stars allowed us to constrain the metallicity distribution across the LMC and its star formation rate (SFR). The LMC stellar population is found to be on average 5-6 Gyr old and is consistent with a mean metallicity corresponding to Z=0.006. These values may however be affected by systematic errors in the underlying stellar models, and by the limited exploration of the possible SFR histories. Instead our method should be particularly useful for detecting variations in the mean metallicity and SFR across the LMC disk. There are well defined regions where both the metallicity and the mean-age of the underlying stellar population span the whole range of grid parameters. The C/M ratio discussed in Paper I is a tracer of the metallicity distribution if the underlying stellar population is older than about a few Gyr. A similar study across the Small Magellanic Cloud is given in Paper III of this series.

Published in Astronomy & Astrophysics 448, 77 (2006)

AGB stars in the Magellanic Clouds. III. The rate of star formation across the SMC

M.-R.L. Cioni^{1}, L. Girardi^{2}, P. Marigo^{3} and H.J. Habing^{4}

1SUPA, School of Physics, University of Edinburgh, IfA, Blackford Hill, Edinburgh EH9 3HJ, UK 2Osservatorio Astronomico di Trieste, INAF, via G. B. Tiepolo 11, 34131 Trieste, Italy 3Dipartimento di Astronomia, Universit di Padova, Vicolo dell’Osservatorio 2, 35122 Padova, Italy 4Sterrewacht Leiden, Niels Bohrweg 2, 2333 RA Leiden, The Netherlands

This article compares the Ks magnitude distribution of Small Magellanic Cloud asymptotic giant branch stars obtained from the DENIS and 2MASS data with theoretical distributions. Theoretical Ks magnitude distributions have been constructed using up-to-date stellar evolution calculations for low and intermediate-mass stars, and in particular for thermally pulsing asymptotic giant branch stars. Separate fits of the magnitude distributions of carbon- and oxygenrich stars allowed us to constrain the metallicity distribution across the galaxy and its star formation rate. The Small Magellanic Cloud stellar population is found to be on average 7-9 Gyr old but older stars are present at its periphery and younger stars are present in the direction of the companion galaxy the Large Magellanic Cloud. The metallicity distribution traces a ring-like structure that is more metal rich than the inner region of the galaxy. The C/M ratio discussed in Paper I is a tracer of the metallicity distribution only if the underlying stellar population is of intermediate-age.

Accepted for publication in Astronomy & Astrophysics Available from astro-ph/0602483

Chandra LETG Observations of Supernova Remnant 1987A

Svetozar A. Zhekov^{1,4}, Richard McCray^{1}, Kazimierz J. Borkowski^{2}, David N. Burrows^{3} and Sangwook Park^{3}

1JILA, University of Colorado, Boulder, CO, USA 2Department of Physics, NCSU, Raleigh, NC, USA 3Department of Astronomy and Astrophysics, Pennsylvania State University, USA 4Current address: Space Research Institute, Sofia, Bulgaria

We discuss the results from deep Chandra LETG observations of the supernova remnant 1987A (SNR 1987A). We find that a distribution of shocks, spanning the same range of velocities (from ∼ 300 to 1700 km s^{−1}) as deduced in the first part of our analysis (Zhekov et al. 2005, ApJL, 628, L127), can account for the entire X-ray spectrum of this object. The post-shock temperature distribution is bimodal, peaking at kT∼ 0.5 and ∼ 3 keV. Abundances inferred from the X-ray spectrum have values similar to those for the inner circumstellar ring, except that the abundances of nitrogen and oxygen are approximately a factor of two lower than those inferred from the optical/UV spectrum. The velocity of the X-ray emitting plasma has decreased since 1999, apparently because the blast wave has entered the main body of the inner circumstellar ring.

Accepted for publication in ApJ Available from astro-ph/0603305

ESO-VLT and Spitzer spectroscopy of IRAS 05328−6827: a massive young stellar object in the Large Magellanic Cloud

Jacco Th. van Loon^{1}, J.M. Oliveira^{1}, P.R. Wood^{2}, A.A. Zijlstra^{3}, G.C. Sloan^{4}, M. Matsuura^{3}, P.A. Whitelock^{5}, M.A.T. Groenewegen^{6}, J.A.D.L. Blommaert^{6}, M.-R.L. Cioni^{7}, S. Hony^{6}, C. Loup^{8} and L.B.F.M. Waters^{9,6}

1Astrophysics Group, School of Physical & Geographical Sciences, Keele University, Staffordshire ST5 5BG, UK 2Research School of Astronomy and Astrophysics, Australian National University, Cotter Road, Weston Creek, ACT 2611, Australia 3School of Physics and Astronomy, University of Manchester, Sackville Street, P.O.Box 88, Manchester M60 1QD, UK 4Department of Astronomy, Cornell University, 108 Space Sciences Building, Ithaca NY 14853-6801, USA 5South African Astronomical Observatory, P.O. Box 9, 7935 Observatory, South Africa 6Instituut voor Sterrenkunde, Celestijnenlaan 200B, B-3001 Leuven, Belgium 7Institute for Astronomy, University of Edinburgh, Royal Observatory, Blackford Hill, Edinburgh EH9 3HJ, UK 8Institut d’Astrophysique de Paris, CNRS, 98bis Boulevard Arago, 75014 Paris, France 9Astronomical Institute, University of Amsterdam, Kruislaan 403, 1098 SJ Amsterdam, The Netherlands

We present the first thermal-infrared spectra of an extra-galactic Young Stellar Object (YSO), IRAS 05328−6827 in the H ii region LHA 120-N 148 in the Large Magellanic Cloud. The observed and modelled spectral energy distribution reveals a massive YSO, M ∼ 20 M_{⊙}, which is heavily-embedded and probably still accreting. The reduced dust content as a consequence of the lower metallicity of the LMC allows a unique view into this object, and together with a high C/O ratio may be responsible for the observed low abundance of water ice and relatively high abundances of methanol and CO_{2} ices.

Published in MNRAS, 364, L71 (2005) Available from astro-ph/0509695

Dust-enshrouded giants in clusters in the Magellanic Clouds

Jacco Th. van Loon^{1}, Jonathan R. Marshall^{1} and Albert A. Zijlstra^{2}

1Astrophysics Group, School of Physical & Geographical Sciences, Keele University, Staffordshire ST5 5BG, UK 2School of Physics and Astronomy, University of Manchester, Sackville Street, P.O.Box 88, Manchester M60 1QD, UK

We present the results of an investigation of post-Main Sequence mass loss from stars in clusters in the Magellanic Clouds, based around an imaging survey in the L′-band (3.8 μm) performed with the VLT at ESO. The data are complemented with JHK_{s} (ESO and 2MASS) and mid-IR photometry (TIMMI2 at ESO, ISOCAM on-board ISO, and data from IRAS and MSX). The goal is to determine the influence of initial metallicity and initial mass on the mass loss and evolution during the latest stages of stellar evolution. Dust-enshrouded giants are identified by their reddened near-IR colours and thermal-IR dust excess emission. Most of these objects are Asymptotic Giant Branch (AGB) carbon stars in intermediate-age clusters, with progenitor masses between 1.3 and ∼5 M_{⊙}. Red supergiants with circumstellar dust envelopes are found in young clusters, and have progenitor masses between 13 and 20 M_{⊙}. Post-AGB objects (e.g., Planetary Nebulae) and massive stars with detached envelopes and/or hot central stars are found in several clusters. We model the spectral energy distributions of the cluster IR objects, in order to estimate their bolometric luminosities and mass-loss rates. The IR objects are the most luminous cluster objects, and have luminosities as expected for their initial mass and metallicity. They experience mass-loss rates in the range from a few 10^{−6} up to 10^{−4} M_{⊙} yr^{−1} (or more), with most of the spread being due to evolutionary effects and only a weak dependence on progenitor mass and/or initial metallicity. About half of the mass lost by 1.3–3 M_{⊙} stars is shed during the superwind phase, which lasts of order 10^{5} yr. Objects with detached shells are found to have experienced the highest mass-loss rates, and are therefore interpreted as post-superwind objects. We also propose a simple method to measure the cluster mass from L′-band images.

Published in A&A, 442, 597 (2005) Available from astro-ph/0507571

An empirical formula for the mass-loss rates of dust-enshrouded red supergiants and oxygen-rich Asymptotic Giant Branch stars

Jacco Th. van Loon^{1}, Maria-Rosa L. Cioni^{2,3}, Albert A. Zijlstra^{4} and Cecile Loup^{5}

1Astrophysics Group, School of Physical & Geographical Sciences, Keele University, Staffordshire ST5 5BG, UK 2European Southern Observatory, Karl-Schwarzschild Straße 2, D-85748 Garching bei M 3Institute for Astronomy, University of Edinburgh, Royal Observatory, Blackford Hill, Edinburgh EH9 3HJ, UK 4School of Physics and Astronomy, University of Manchester, Sackville Street, P.O.Box 88, Manchester M60 1QD, UK 5Institut d’Astrophysique de Paris, 98bis Boulevard Arago, F-75014 Paris, France

We present an empirical determination of the mass-loss rate as a function of stellar luminosity and effective temperature, for oxygen-rich dust-enshrouded Asymptotic Giant Branch stars and red supergiants. To this aim we obtained optical spectra of a sample of dust-enshrouded red giants in the Large Magellanic Cloud, which we complemented with spectroscopic and infrared photometric data from the literature. Two of these turned out to be hot emission-line stars, of which one is a definite B[e] star. The mass-loss rates were measured through modelling of the spectral energy distributions. We thus obtain the mass-loss rate formula log Ṁ = −5.65+1.05 log(L/10, 000 L_{⊙})−6.3 log(T_{eff}/3500 K), valid for dust-enshrouded red supergiants and oxygen-rich AGB stars. Despite the low metallicity of the LMC, both AGB stars and red supergiants are found at late spectral types. A comparison with galactic AGB stars and red supergiants shows excellent agreement between the mass-loss rate as predicted by our formula and that derived from the 60 μm flux density for dust-enshrouded objects, but not for optically bright objects. We discuss the possible implications of this for the mass-loss mechanism.

Published in A&A, 438, 273 (2005) Available from astro-ph/0504379

Very Large Telescope three micron spectra of dust-enshrouded red giants in the Large Magellanic Cloud

Jacco Th. van Loon^{1}, Jonathan R. Marshall^{1}, Martin Cohen^{2}, Mikako Matsuura^{3}, Peter R. Wood^{4}, Issei Yamamura^{5} and Albert A. Zijlstra^{3}

1Astrophysics Group, School of Physical & Geographical Sciences, Keele University, Staffordshire ST5 5BG, UK 2Radio Astronomy Lab, 601 Campbell Hall, University of California at Berkeley, Berkeley CA 94720-3411, USA 3Department of Physics and Astronomy, University of Manchester, Sackville Street, P.O.Box 88, Manchester M60 1QD, UK 4Research School of Astronomy and Astrophysics, Australian National University, Cotter Road, Weston Creek, ACT 2611, Australia 5Institute of Space and Astronautical Science, 3-1-1 Yoshinodai, Sagamihara, Kanagawa 229-8510, Japan

We present ESO/VLT spectra in the 2.9–4.1 μm range for a large sample of infrared stars in the Large Magellanic Cloud (LMC), selected on the basis of MSX and 2MASS colours to be extremely dust-enshrouded AGB star candidates. Out of 30 targets, 28 are positively identified as carbon stars, significantly adding to the known population of optically invisible carbon stars in the LMC. We also present spectra for six IR-bright stars in or near three clusters in the LMC, identifying four of them as carbon stars and two as oxygen-rich supergiants. We analyse the molecular bands of C_{2}H_{2} at 3.1 and 3.8 μm, HCN at 3.57 μm, and sharp absorption features in the 3.70–3.78 μm region that we attribute to C_{2}H_{2}. There is evidence for a generally high abundance of C_{2}H_{2} in LMC carbon stars, suggestive of high carbon-to-oxygen abundance ratios at the low metallicity in the LMC. The low initial metallicity is also likely to have resulted in less abundant HCN and CS. The sample of IR carbon stars exhibits a range in C_{2}H_{2}:HCN abundance ratio. We do not find strong correlations between the properties of the molecular atmosphere and circumstellar dust envelope, but the observed differences in the strengths and shapes of the absorption bands can be explained by differences in excitation temperature. High mass-loss rates and strong pulsation would then be seen to be associated with a large scale height of the molecular atmosphere.

Published in A&A, 447,971 (2006) Available from astro-ph/0510510

VLT/UVES Observations of Interstellar Molecules and Diffuse Bands in the Magellanic Clouds

D. E. Welty^{1}, S. R. Federman^{2}, R. Gredel^{3}, J. A. Thorburn^{1} and D. L. Lambert^{4}

1Univ. of Chicago 2Univ. of Toledo 3Max-Planck-Inst. für Astronomie 4Univ. of Texas

We discuss the abundances of interstellar CH, CH^{+}, and CN in the Magellanic Clouds, derived from spectra of 7 SMC and 13 LMC stars obtained (mostly) with the VLT/UVES. CH and/or CH^{+} have now been detected toward 3 SMC and 9 LMC stars; CN is detected toward Sk 143 (SMC) and Sk−67 2 (LMC). These data represent nearly all the optical detections of these molecular species in interstellar media beyond the Milky Way. In the LMC, the CH/H_{2} ratio is comparable to that found for diffuse Galactic molecular clouds in four sight lines, but is lower by factors of 2.5–4.0 in two others. In the SMC, the CH/H_{2} ratio is comparable to the local Galactic value in one sight line, but is lower by factors of 10–15 in two others. The abundance of CH in the Magellanic Clouds thus appears to depend on local physical conditions — and not just on metallicity. In both the SMC and LMC, the observed relationships between the column density of CH and those of CN, CH^{+}, Na i, and K i are generally consistent with the trends observed in our Galaxy.

Using existing data for the rotational populations of H_{2} in these sight lines, we estimate temperatures, radiation field strengths, and local hydrogen densities for the diffuse molecular gas. The inferred temperatures range from about 45 to 90 K, the radiation fields range from about 1 to 900 times the typical local Galactic field, and the densities (in most cases) lie between 100 and 600 cm^{−3}. Densities estimated from the observed N(CH), under the assumption that CH is produced via steady-state gas-phase reactions, are considerably higher than those derived from H_{2}. Much better agreement is found by assuming that the CH is made via the (still undetermined) process(es) responsible for the observed CH^{+}. A significant fraction of the CH and CH^{+} in diffuse molecular material in the SMC and LMC may be produced in photon-dominated regions. The excitation temperature obtained from the populations of the two lowest CN rotational levels toward Sk−67 2 is quite consistent with the temperature of the cosmic microwave background radiation measured with COBE.

Toward most of our targets, the UVES spectra also reveal absorption at velocities corresponding to the Magellanic Clouds ISM from several of the strongest of the diffuse interstellar bands (at 5780, 5797, and 6284 Å). On average, the three DIBs are weaker by factors of 7 to 9 (LMC) and about 20 (SMC), compared to those typically observed in Galactic sight lines with similar N(H i) — presumably due to the lower metallicities and stronger radiation fields in the LMC and SMC. The three DIBs are also weaker (on average, but with some exceptions) by factors of order 2 to 6, relative to E(B − V), N(Na i), and N(K i) in the Magellanic Clouds. The detection of several of the so-called “C_{2} DIBs” toward Sk 143 and Sk−67 2, with strengths similar to those in comparable Galactic sight lines, however, indicates that no single, uniform scaling factor (e.g., one related to metallicity) applies to all DIBs (or for all sight lines) in the Magellanic Clouds.

Accepted for publication in ApJS Available from astro-ph/0603332

On the Unusual Depletions toward Sk 155, or What Are the Small Magellanic Cloud Dust Grains Made of?

Aigen Li^{1}, K.A. Misselt^{2} and Y.J. Wang^{3}

1Department of Physics and Astronomy, University of Missouri, Columbia, MO 65211 2Steward Observatory, University of Arizona, Tucson, AZ 85721 3Department of Physics, Hunan Normal University, Changsha 410071, P.R. China

The dust in the Small Magellanic Cloud (SMC), an ideal analog of primordial galaxies at high redshifts, differs markedly from that in the Milky Way by exhibiting a steeply rising far-ultraviolet extinction curve, an absence of the 2175 Å extinction feature, and a local minimum at ∼12 μm in its infrared emission spectrum, suggesting the lack of ultrasmall carbonaceous grains (i.e. polycyclic aromatic hydrocarbon molecules) which are ubiquitously seen in the Milky Way. While current models for the SMC dust all rely heavily on silicates, recent observations of the SMC sightline toward Sk 155 indicated that Si and Mg are essentially undepleted and the depletions of Fe range from mild to severe, suggesting that metallic grains and/or iron oxides, instead of silicates, may dominate the SMC dust. However, in this Letter we apply the Kramers-Kronig relation to demonstrate that neither metallic grains nor iron oxides are capable of accounting for the observed extinction; silicates remain as an important contributor to the extinction, consistent with current models for the SMC dust.

Published in The Astrophysical Journal Letters, vol. 640, L151-L154 (the April 01, 2006 issue) Available from astro-ph/0603573 and from http://www.missouri.edu/∼lia/

The Effective Temperatures and Physical Properties of Magellanic Cloud Red Supergiants: The Effects of Metallicity

Emily M. Levesque^{1,2}, Philip Massey^{1}, K. A. G Olsen^{3}, Bertrand Plez^{4}, Georges Meynet^{5} and Andre Maeder^{5}

1Lowell Observatory 2MIT 3CTIO/NOAO 4GRAAL CNRS, Univ. de Montpellier II

We present moderate-resolution optical spectrophotometry of 36 red supergiants (RSGs) in the LMC and 39 RSGs in the SMC. Using the MARCS stellar atmosphere models to fit this spectrophotometry, we determine the reddenings, effective temperatures and other physical properties, such as bolometric luminosity and effective stellar radii, and compare these to stellar evolutionary models. As a self-consistency check, we also compare the broad-band colors (V − K)_{0} and (V − R)_{0} with the models. The (V − R)_{0} results are in good agreement with those from fitting the optical spectrophotometry, but the (V − K)_{0} results show metallicity-dependent systematic differences, amounting to 3-4% in effective temperature, and 0.2 mag in bolometric luminosity, at the metallicity of the SMC; we conclude that this is likely due to the limitations of static 1D models, as spectra of RSGs in the optical and IR may reflect different atmospheric conditions due to the large surface granulation present in these stars. We adopt the scales indicated by the optical spectrophotometry and (V − R)_{0} colors, but accept that there is still some uncertainty in the absolute temperature scales. We find that the effective temperature scales for the LMC and SMC K-type supergiants agree with each other and with that of the Milky Way, while for M-type supergiants the scales are cooler than the Galactic scale by 50 K and 150 K, respectively. This is in the sense that one would expect: since the spectral classification of RSGs is based on the line strengths of TiO, stars with lower abundances of these elements have to be cooler in order to have the same strength. However, this effect is not sufficient to explain the shift in average RSG spectral type between the three galaxies. Instead, it is the effect that metallicity has on the coolest extent of the evolution of a star that is primarily responsible. Our new results bring the RSGs into much better agreement with stellar evolutionary theory, although the SMC RSGs show a considerably larger spread in effective temperatures at a given luminosity than do the LMC stars. This is expected due to the larger effects of rotational mixing in lower-metallicity stars, as higher helium abundance at the surface would lead to higher effective temperatures in the RSG phase. We also find that the distribution of reddening of RSGs in the Clouds is skewed significantly towards higher values, consistent with our recent finding that Galactic RSGs show extra extinction due to circumstellar dust.

Accepted for publication in ApJ Available from astro-ph/0603596 and from http://www.lowell.edu/users/massey/MCRSG.pdf.gz

The Hubble Constant from Type Ia Supernova Calibrated with the Linear and Non-Linear Cepheid Period-Luminosity Relation

C. Ngeow^{1} and S. Kanbur^{2}

1UIUC 2SUNY-Oswego

It is well-known that the peak brightness of the Type Ia supernovae calibrated with Cepheid distances can be used to determine the Hubble constant. The Cepheid distances to host galaxies of the calibrating supernovae are usually obtained using the period-luminosity (PL) relation derived from Large Magellanic Cloud (LMC) Cepheids. However recent empirical studies provide evidence that the LMC PL relation is not linear. In this Letter we determine the Hubble constant using both the linear and non-linear LMC Cepheid PL relations as calibrating relations to four galaxies that hosted Type Ia supernovae. Our results suggest that the obtained values of the Hubble constant are similar. However a typical error of ∼ 0.03 mag. has to be added (in quadrature) to the systematic error for the Hubble constant when the linear LMC PL relation is used, assuming that the LMC PL relation is indeed non-linear. This is important in minimizing the total error of the Hubble constant in the era of precision cosmology. The Hubble constants calibrated from the linear and non-linear LMC PL relation are H_{0} = 74.92±2.28 (random) ±5.06 (systematic) km/s/Mpc and H_{0} = 74.37 ± 2.27 (random) ±4.92 (systematic) km/s/Mpc, respectively. Hubble constants calculated using the Galactic PL relations are also briefly discussed and presented in the last section of this Letter.

Accepted for publication in ApJL Available from astro-ph/0603643

Spitzer Space Telescope detection of the young supernova remnant 1E 0102.2-7219

Snežana Stanimirović^{1}, Alberto D. Bolatto^{1}, Karin Sandstrom^{1}, Adam K. Leroy^{1}, Joshua D. Simon^{1,2}, B. M. Gaensler^{3}, Ronak Y. Shah^{4} and James M. Jackson^{4}

1Radio Astronomy Lab, UC Berkeley, 601 Campbell Hall, Berkeley, CA 94720 2Department of Astronomy, California Institute of Technology, 1200 E. California Blvd, MS 105-24, Pasadena, CA 91125 3Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138 4Institute for Astrophysical Research, Boston University, 725 Commonwealth Avenue, Boston, MA 02215

We present infrared observations of the young, oxygen-rich supernova remnant 1E 0102.2-7219 (E0102) in the Small Magellanic Cloud, obtained with the Spitzer Space Telescope. The remnant is detected at 24 μm but not at 8 or 70 μm and has a filled morphology with two prominent filaments. We find evidence for the existence of up to 8 × 10^{−4} M_{⊙} of hot dust (T_{d} ∼ 120 K) associated with the remnant. Most of the hot dust is located in the central region of E0102 which appears significantly enhanced in infrared and radio continuum emission relative to the X–ray emission. Even if all of the hot dust was formed in the explosion of E0102, the estimated mass of dust is at least 100 times lower that what is predicted by some recent theoretical models.

Published in ApJ, 632, L103 (2005)

Mid-infrared spectroscopy of carbon stars in the Small Magellanic Cloud

Sloan, G.C.^{1}, Kraemer, K.E.^{2}, Matsuura, M.^{3,4}, Wood, P.R.^{5}, Price, S.D.^{2} and Egan, M.P.^{2}

1Cornell University 2Air Force Research Laboratory 3University of Manchester 4Queen’s University of Belfast 5Australian National University

We have observed a sample of 36 objects in the Small Magellanic Cloud (SMC) with the Infrared Spectrometer on the Spitzer Space Telescope. Nineteen of these sources are carbon stars. An examination of the near- and midinfrared photometry shows that the carbon-rich and oxygen-rich dust sources follow two easily separated sequences. A comparison of the spectra of the 19 carbon stars in the SMC to spectra from the Infrared Space Observatory (ISO) of carbon stars in the Galaxy reveals significant differences. The absorption bands at 7.5 μm and 13.7 μm due to C_{2}H_{2} are stronger in the SMC sample, and the SiC dust emission feature at 11.3 μm is weaker. Our measurements of the MgS dust emission feature at 26–30 μm are less conclusive, but this feature appears to be weaker in the SMC sample as well. All of these results are consistent with the lower metallicity in the SMC. The lower abundance of SiC grains in the SMC may result in less efficient carbon-rich dust production, which could explain the excess C_{2}H_{2} gas seen in the spectra. The sources in the SMC with the strongest SiC dust emission tend to have redder infrared colors than the other sources in the sample, which implies more amorphous carbon, and they also tend to show stronger MgS dust emission. The weakest SiC emission features tend to be shifted to the blue; these spectra may arise from low-density shells with large SiC grains.

Accepted for publication in ApJ, 645 Available from astro-ph/0603607 and from http://isc.astro.cornell.edu/∼sloan/library/smcc.html

Coronal emission from the shocked circumstellar ring of SN 1987A

Per Gröningsson^{1}, Claes Fransson^{1}, Peter Lundqvist^{1}, Tanja Nymark^{1}, Natalia Lundqvist^{1}, Roger Chevalier^{2}, Bruno Leibundgut^{3} and Jason Spyromilio^{3}

1Department of Astronomy, Stockholm University, Sweden 2Department of Astronomy, University of Virginia 3European Southern Observatory, Garching, Germany

High resolution spectra with UVES/VLT of SN 1987A from December 2000 until November 2005 show a number of high ionization lines from gas with velocities of ∼ ±350 km s^{−1}, emerging from the shocked gas formed by the ejecta – ring collision. These include coronal lines from [Fe x], [Fe xi] and [Fe xiv] which have increased by a factor of ∼ 20 during the observed period. The evolution of the lines is similar to that of the soft X-rays, indicating that they arise in the same component. The line ratios are consistent with those expected from radiative shocks with velocity 310 − 390 km s^{−1}, corresponding to a shock temperature of (1.6 − 2.5) 10^{6} K. A fraction of the coronal emission may, however, originate in higher velocity adiabatic shocks.

Submitted to Astronomy & Astrophysics Available from astro-ph/0603815

Conference Papers

Old Main-Sequence Turnoff Photometry in the SMC

Noelia E. D. Noel^{1}, Carme Gallart^{1}, Edgardo Costa^{2} and René A. Méndez^{2}

1Instituto de Astrofisica de Canarias 2Departamento de Astronomia, Universidad de Chile

We present ground-based B and R-band color-magnitude diagrams (CMDs) of unprecedented depth for twelve fields in the Small Magellanic Cloud (SMC). They reach the oldest main-sequence turnoffs and cover a wide range of galactocentric distances up to ∼4^{∘} from the SMC center, and are located at different position angles. A picture of the stellar content in our SMC fields is presented, through the comparison with theoretical isochrones. Our study confirms the existence of strong population gradients and spatial variation in the SMC stellar content. None of the SMC fields presented here are dominated by old stellar populations which proves that at ∼4^{∘} from the SMC center we do not reach an old stellar halo similar to that of the Milky Way.

Oral contribution, published in To appear in Revista Mexicana de Astronomía y Astrofísica, Serie de Conferencias, LARIM 2005 Available from astro-ph/0603319 and from ftp://ftp.iac.es/out/noelia/

Spitzer IRS spectra of Luminous 8 μm Sources in the Large Magellanic Cloud

Catherine L. Buchanan^{1}, Joel H. Kastner^{1}, William J. Forrest^{2}, Bruce J. Hrivnak^{3}, Raghvendra Sahai^{4}, Michael Egan^{5}, Adam Frank^{2} and Cecilia Barnbaum^{6}

1Center for Imaging Science, Rochester Institute of Technology 2Department of Physics & Astronomy, University of Rochester 3Department of Physics and Astronomy, Valparaiso University 4NASA/JPL 5Air Force Research Laboratory; OASD (NII) Space Programs 6Valdosta University

We have produced an atlas of Spitzer Infrared Spectrograph (IRS) spectra of mass-losing, evolved stars in the Large Magellanic Cloud. These stars were selected to have high mass-loss rates and so contribute significantly to the return of processed materials to the ISM. Our high-quality spectra enable the determination of the chemistry of the circumstellar envelope from the mid-IR spectral features and continuum. We have classified the spectral types of the stars and show that the spectral types separate clearly in infrared color-color diagrams constructed from 2MASS data and synthetic IRAC/MIPS fluxes derived from our IRS spectra. We present diagnostics to identify and classify evolved stars in nearby galaxies with high confidence levels using Spitzer and 2MASS photometry. Comparison of the spectral classes determined using IRS data with the IR types assigned based on NIR colors also revealed a significant number of misclassifications and enabled us to refine the NIR color criteria resulting in more accurate NIR color classifications of dust-enshrouded objects.

Poster contribution, published in ASP conf. ser., ”IR Diagnostics of Galaxy Evolution”, 14-16 November 2005, Pasadena Available from astro-ph/0603443

The Far-IR Radio Continuum Correlation in the Small Magellanic Cloud

Karin Sandstrom^{1}, Alberto Bolatto^{1}, Adam Leroy^{1}, Snežana Stanimirović^{1}, Joshua Simon^{2}, Lister Staveley-Smith^{3} and Ronak Shah^{4}

1Radio Astronomy Lab, UC Berkeley 2Department of Astronomy, California Institute of Technology 3Australia Telescope National Facility, CSIRO 4Institute for Astrophysical Research, Boston University

We present the results of a study of the far-IR radio continuum correlation within the Small Magellanic Cloud using maps of the far-infrared luminosity derived from MIPS 70 and 160 μm mosaics and radio continuum maps from a combined ATCA-Parkes survey. With these data, we can study the correlation within the SMC down to scales of about 35 parsecs. We present preliminary evidence that the correlation holds even at these small scales and that the slope of the correlation in the SMC is similar, though slightly steeper, than that found in previous studies of larger galaxies.

Poster contribution, published in ASP Conference Series: Infrared Diagnostics Of Galaxy Evolution Available from http://celestial.berkeley.edu/spitzer/publications/sandstrom irevolve.pdf

The young SNR 1E 0102.2-7219: testing dust formation in primordial galaxies

Snežana Stanimirović^{1}, Alberto D. Bolatto^{1}, Karin Sandstrom^{1}, Adam Leroy^{1}, Joshua D. Simon^{2}, B. M. Gaensler^{3}, Ronak Shah^{4} and James M. Jackson^{4}

1Radio Astronomy Lab, UC Berkeley, 601 Campbell Hall, Berkeley, CA 94720 2Department of Astronomy, California Institute of Technology, 1200 E. California Blvd, MS 105-24, Pasadena, CA 91125 3Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138 4Institute for Astrophysical Research, Boston University, 725 Commonwealth Avenue, Boston, MA 02215

We present infrared (IR) observations of the young, oxygen-rich supernova remnant 1E 0102.2-7219 (E0102) in the Small Magellanic Cloud (SMC), obtained with the Spitzer Space Telescope. The remnant is detected only at 24 μm and has a filled morphology with two prominent filaments. We find evidence for the existence of up to 8 × 10^{−4} M_{⊙} of hot dust (T_{d} ∼ 120 K) associated with the remnant. Even if all of the hot dust was formed in the explosion of E0102, the estimated mass of dust is at least 100 times lower that what is predicted by some recent theoretical models. The implied dust formation efficiency in the SMC is low, < 15%, and challenges theoretical models for dust formation in primordial galaxies. Most of the hot dust in E0102 is located in the central region which appears significantly enhanced in IR and radio continuum emission relative to the X–ray emission. Curiously, we find a compact source of excess IR emission close to E0102’s center. This could be a synchrotron jet encountering a dusty interstellar medium, or alternatively, a dust condensate embedded in a cooler interstellar cloud.

Oral contribution, published in The Spitzer Science Center 2005 Conference: ”Infrared Diagnostics of Galaxy Evolution” Available from http://celestial.berkeley.edu/spitzer/publications/

Infrared Sources in the Small Magellanic Cloud: First Results

Joshua D. Simon^{1}, Alberto D. Bolatto^{2}, Snežana Stanimirović^{2}, Ronak Y. Shah^{3}, Adam Leroy^{2} and Karin Sandstrom^{2}

1Caltech 2UC Berkeley 3Boston University

We have imaged the entire Small Magellanic Cloud (SMC), one of the two nearest star-forming dwarf galaxies, in all seven IRAC and MIPS bands. The low mass and low metallicity (1/6 solar) of the SMC make it the best local analog for primitive galaxies at high redshift. By studying the properties of dust and star formation in the SMC at high resolution, we can gain understanding of similar distant galaxies that can only be observed in much less detail.

In this contribution, we present a preliminary analysis of the properties of point sources detected in the Spitzer Survey of the Small Magellanic Cloud (S^{3}MC). We find ∼ 400, 000 unresolved or marginally resolved sources in our IRAC images, and our MIPS 24 micron mosaic contains ∼ 17, 000 point sources. Source counts decline rapidly at the longer MIPS wavelengths. We use color-color and color-magnitude diagrams to investigate the nature of these objects, cross-correlate their positions with those of known sources at other wavelengths, and show examples of how these data can be used to identify interesting classes of objects such as carbon stars and young stellar objects. For additional examples of some of the questions that can be studied with these data, please see the accompanying contributions by Alberto Bolatto (survey information and images), Adam Leroy (dust and gas in a low-metallicity environment), Karin Sandstrom (far infrared-radio continuum correlation), and Snežana Stanimirović (on a young supernova remnant in the SMC). The mosaic images and point source catalogs we have made have been released to the public on our website (http://celestial.berkeley.edu/spitzer).

Poster contribution, published in Infrared Diagnostics of Galaxy Evolution Available from astro-ph/0603834 and from http://celestial.berkeley.edu/spitzer/publications/simon irevolve.pdf

Thesis

Mass loss from dust-enshrouded Asymptotic Giant Branch stars and red supergiants in the Large Magellanic Cloud

Jonathan Marshall^{1}

1Astrophysics Group, School of Physical & Geographical Sciences, Keele University, Staffordshire ST5 5BG, UK. jrm@astro.keele.ac.uk

The process of mass-loss from evolved stars is the single largest contributor of matter back into the ISM. Intense mass-loss during the AGB phase of low-intermediate mass stars via a radiatively-driven wind can lead the stars to become enshrouded in an optically-thick layer of dust which condenses out of an extended molecular atmosphere. This thesis attempts to gain further insights into the mass-loss process that is presently poorly understood.

We used the Parkes radio telescope to observe dust-enshrouded AGB stars and supergiants in the LMC and SMC, deriving the speed of the superwind from the double-peaked OH maser profiles. Out of 8 targets in the LMC we detected 5, of which 3 are new detections. Our results confirm the simple theory for radiatively driven winds, this verifies the scaling relations we use in determining mass-loss rates and allows us to speculate on the chemical enrichment at different metallicities.

From investigating mass-loss from clusters in the Magellanic Clouds we find that the mass-loss rate increases with larger progenitor mass, possibly due to a dependence on the initial metallicity or the stellar luminosity. We investigate the dust-enshrouded carbon star LI-LMC 1813 in more depth and derive an accurate mass-loss rate and the stellar parameters, mass and metallicity. It is now one of the few AGB stars currently undergoing the superwind phase for which values for the fundamental astrophysical parameters are known.

With the ESO Very Large Telescope we obtained 3–4 μm spectra of IR stars in the LMC. 28 of 30 targets are identified as carbon stars, significantly adding to the known population of optically invisible carbon stars in the LMC. We find evidence for a high abundance of C_{2}H_{2}, suggestive of high carbon-to-oxygen abundance ratios at the low metallicity which would explain the large population of carbon stars.

Defended on 14th December 2005. Thesis supervisor: Dr. J.Th. van Loon (Keele University, UK)